OpenStudy (anonymous):

I have a problem where I need to figure out how many pounds of fruit an 18 ft & a 20 tree will bear. Here is the formula: p-24=31/11(h-5). For the life of me I am stuck and can't figure it out. Please help and show me.

OpenStudy (anonymous):

do you not just plug in 18 for the height and solve for p?

OpenStudy (anonymous):

I did that ans when I submitted the answer it said it was wrong

OpenStudy (anonymous):

That is why I am asking for help I am not sure why and what I need to do

OpenStudy (anonymous):

is the formula p-24 = 31/(11(h-5))

OpenStudy (anonymous):

I used the formula I just put and got 24.94 pounds for the 18 ft tree and 24.56 pounds for the 20 ft tree.

OpenStudy (anonymous):

no Here is the ordered pair (5,24) & (16,55) then I did 55-24/16-5 = 31/11

OpenStudy (anonymous):

the equation came to p-24=31/11(h-5)

OpenStudy (anonymous):

I am suppose to take that equation and predict the # of lbs of fruit a 18 & a 20 ft tree will produce

OpenStudy (anonymous):

dang. I left the page and when I went back the problem was gone so now I will have to do a different one. Sorry I will have to post a different problem.

OpenStudy (anonymous):

oh you are trying to find the equation of the line? y = mx +b m = (55-24)/(16-5) = 31/11 y = 31/11(x - 5) + 24 for x=18 y = (31/11)*3+24 = 32.45 for x = 20 y = (31/11)*5+24 = 38.09

OpenStudy (anonymous):

you line equation was incorrect. Use point-slope form. \[y = m(x -x _{1})+ y_{1}\]

OpenStudy (anonymous):

that is not the formula they are using. They are using m= p2-p1/h2-h1=

OpenStudy (anonymous):

yeah that is the slope equation. Mine is the slope equation too...just rearranged.

OpenStudy (anonymous):

ok i was just following the examples they give

OpenStudy (anonymous):

I can rearrange yours to do the same thing: p2 - p1 = m(h2-h1) p2 = m(h2-h1) + p1

OpenStudy (anonymous):

where h2 and p2 would be your variables

OpenStudy (anonymous):

and h1 and p1 would be a known point

OpenStudy (anonymous):

ok the new formula, since I had to change problems, is p-18=20/3(h-6)

OpenStudy (anonymous):

p1 & p2 = 18 & 58 and h1 & h2 = 6 & 12 respectively

OpenStudy (anonymous):

how do I simplify that equation?

OpenStudy (anonymous):

p = (20/3)*(h-6)+18

OpenStudy (anonymous):

ok so I am substituting 18 for h

OpenStudy (anonymous):

I got 132

OpenStudy (anonymous):

ok so p = (20/3)*(18-6) + 18 p = (20/3)*12 + 18 = 98

OpenStudy (anonymous):

I see I forgot to do the () first

OpenStudy (anonymous):

ok. Did you get 98?

OpenStudy (anonymous):

yes I did

OpenStudy (anonymous):

That looks like an easier formula to use

OpenStudy (anonymous):

Ok so that is the first answer. What is the other height value?

OpenStudy (anonymous):

20 ft

OpenStudy (anonymous):

Yeah it is the same as the one they are giving you. It is just rearranged algebraically.

OpenStudy (anonymous):

Thank you. To mee it looks easier to use as long as I remember to do the ( ) first.

OpenStudy (anonymous):

ok so use the same formula. Tell me what you get.

OpenStudy (anonymous):

334/3

OpenStudy (anonymous):

it said the 132 was wrong

OpenStudy (anonymous):

132 was wrong. Remember we said the answer was 98. 132 was the answer you got when you didn't do the () first

OpenStudy (anonymous):

334/3 is the right answer for the height of 20

OpenStudy (anonymous):

ok forgot to write that down. i will try again

OpenStudy (anonymous):

ok 98 was right sorry did I get ur heart going?

OpenStudy (anonymous):

Haha. No I was pretty sure that 98 was right

OpenStudy (anonymous):

but it says to round it to the nearest whole nmuber

OpenStudy (anonymous):

ok then the answer would be 111

OpenStudy (anonymous):

334/3 = 111.33 which is approximated to 111 when rounding to the nearest whole number

OpenStudy (anonymous):

yep right i just though about that too.

OpenStudy (anonymous):

ok last thing. i need to predict the height of a tree that produces 88 lbs. can you tell me how to flip the formula around to solve for that?

OpenStudy (anonymous):

so the formula we were using was: p = (20/3)*(h-6) + 18 subtract 18 from each side p-18 = (20/3)*(h-6) divide each side by (20/3) this is the same as multiplying by 3/20 (3/20)*(p-18) = h-6 now add 6 to each side our new equation is h = (3/20)*(p-18) + 6

OpenStudy (anonymous):

so plug in 88 and tell me what you get

OpenStudy (anonymous):

I got h=16.5

OpenStudy (anonymous):

ya that is what I got

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

rounded to the nearest whole number is 17

OpenStudy (anonymous):

it is telling me to round to the nearest tenth

OpenStudy (anonymous):

then the answer is 16.5

OpenStudy (anonymous):

right. THANK YOU very much. I got another one similiar which i will do and see how it turn out.

OpenStudy (anonymous):

ok good. Glad I could help

OpenStudy (anonymous):

take care